OpenStudy (anonymous):

(12x^2 +19x +5) divided by 3x+1

OpenStudy (gorv):

\[12*x^2+19x+5=12x^2+15x+4x+5=3x(4x+5)+1*(4x+5)=(3x+1)*(4x+5)\]

OpenStudy (gorv):

\[\frac{ (3x+1)*(4x+5) }{ 3x+1 }=4x+5\]

OpenStudy (anonymous):

\[\frac{ 12x ^{2}+4x+15x+5 }{ 3x+1 }=\frac{ 4x(3x+1)+5(3x+1) }{ 3x+1}=4x+5\]

OpenStudy (gorv):

@zanayp

OpenStudy (anonymous):

thank you

OpenStudy (gorv):

wlcme :) @zanayp